There is a 90% chance that a pizza place will make a delivery in less than 30 min. what is the probability that in a sample of 10 pizzas ordered from a certain pizza place at different times of the day atleast 7 will be delivered in less than 30 min?
Answer by stanbon(57967)

There is a 90% chance that a pizza place will make a delivery in less than 30 min. what is the probability that in a sample of 10 pizzas ordered from a certain pizza place at different times of the day at least 7 will be delivered in less than 30 min?
-----------------
Binomial Problem with n = 10 ; p = 0.9
----
P(7<= x <=10) = 1-P(0<= x <=6) = 1 - binomcdf(10,0.9,6)
= 1 - 0.01280
---
= 0.9872
=============
Cheers,
Stan H.